OpenStudy (anonymous):

Prove: (sin pi/3)*(cos pi /6)+(tan pi/6)*(tan pi/3)+cos (pi)=3/4

OpenStudy (anonymous):

sin pi/3 = sqrt(3)/2 cos pi/6= sqrt(3)/2 tan pi/6= 1/sqrt(3) tan pi/3 = sqrt(3) cos pi = -1

OpenStudy (raden):

proof :)

OpenStudy (anonymous):

not proved

OpenStudy (anonymous):

1/sqrt(3) *sqrt(3) "sqrt(3) will cancel out

OpenStudy (anonymous):

3/4-1

OpenStudy (anonymous):

-1/4 :/ ??

OpenStudy (raden):